Three cases are provided below concerning a plant asset currently held for sale. Exercise 12-77 Recording Impairment Loss on Asset Held for Sale LO8 Case Original Cost Selling Costs Accumulated Depreciation 50,000 150,000 Fair Value 2 3 $96,000 45,000 280,000 45,000 30,000 140,000 $1,500 1,000 7,000 1,000 Which cases(s), if any, require an impairment loss to be recognized, and for what amount? Record the entry for the impairment loss on the case(s) identified in part a (if any). a. b.
Answer #1
a) Case 1 Case 2 Case 3
Carrying Value = Cost - Accumulated Depreciation $ 46,000.00 $ 44,000.00 $ 1,30,000.00
Fair value less costs to sell $ 43,500.00 $ 29,000.00 $ 1,33,000.00
Decrease in value (impairment loss recongnized) $   2,500.00 $ 15,000.00 $    -3,000.00

If carrying amount > ‘fair value less costs to sell’ recognise an ‘impairment loss’ (expense) , Case 1 and Case 2 is impairment loss

b) Journal (case 1)
Impairment loss (expens) $   2,500.00
                  Plant: accumulated impairment loss $   2,500.00
Journal (case 2)
Impairment loss (expens) $ 15,000.00
                  Plant: accumulated impairment loss $ 15,000.00
a) Case 1 Case 2 Case 3
Carrying Value = Cost - Accumulated Depreciation =96000-50000 =45000-1000 =280000-150000
Fair value less costs to sell =45000-1500 =30000-1000 =140000-7000
Decrease in value (impairment loss recongnized) =B2-B3 =C2-C3 =D2-D3

If carrying amount > ‘fair value less costs to sell’ recognise an ‘impairment loss’ (expense) , Case 1 and Case 2 is impairment loss

b) Journal (case 1)
Impairment loss (expens) =B4
                  Plant: accumulated impairment loss =B8
Journal (case 2)
Impairment loss (expens) =C4
                  Plant: accumulated impairment loss =B12